OpenStudy (anonymous):

Can I get help? What is the product in lowest terms 1/6 x 5/7

OpenStudy (anonymous):

It's \[\frac{ 1 }{ 6 } * \frac{ 5 }{ 7 } = ?\] In lowest terms

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

multiply numerators multiply denominators

OpenStudy (anonymous):

Ok, let me plug them in real quick

ganeshie8 (ganeshie8):

\[\frac{ 1 }{ 6 } * \frac{ 5 }{ 7 } = \frac{ 1 *5}{ 6*7 } = ? \]

OpenStudy (anonymous):

It's 8.4 as a decimal

ganeshie8 (ganeshie8):

how ? you should leave it in fraction form : \[\frac{ 1 }{ 6 } * \frac{ 5 }{ 7 } = \frac{ 1 *5}{ 6*7 } = \dfrac{5}{42} \]

ganeshie8 (ganeshie8):

we're done !

OpenStudy (anonymous):

Oh, ok, thanks @ganeshie8

ganeshie8 (ganeshie8):

yw !

OpenStudy (anonymous):

hoose all properties that were used to simplify the following problem: • 53 • 53 • • 53 • 1 53 commutative property of addition commutative property of multiplication associative property of addition associative property of multiplication multiplicative identity multiplicative inverse distributive property this is the first problem I need help on I can't seem to find out which one's are the right operations @ganeshie8 can you help with this?

OpenStudy (anonymous):

@Aliypop I asked ganeshie8 too help

OpenStudy (anonymous):

*choose

OpenStudy (anonymous):

@anteater